Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2017-06-02

Night report 2017-06-02

Observers: G. Pignata, L. Wang
TAT on duty: J. Anderson, R. Cartier
Support team: L. Galbany

UT=12.15 Skyflat Arc
Low wind, low humidity, lots of clouds, technical problems with EFOSC starting 4.0 UT
UT/Target/RA/DEC/time spent(min)/slit/grism or filter, description

22:45 LTT3864    10:32:13.90 -35:37:42.4 15min 1.0" Gr11+Gr13+Gr16 seeing 0.7" standard
22:58 problem with the display need to reboot the system
23:37 AT2017awz  11:07:35.48  22:51:04.5 50min 1.0" Gr13 seeing 0.7"
00:20 AT2016iyd  07:37:09.76 -52:19:04.1 15min gri seeing 0.6"
00:35 AT2017ben  09:18:48.44 -12:15:33.1 40min BVgri  seeing 0.6" No SN visible, clouds?
01:15 SN2017dft  10:57:32.05 -14:50:06.0 120min 1.0" gr13 seeing 0.7"
03:15 AT2017eef  11:38:36.85 -09:01:47.9 class 40min as Ia-91T at +16d z=0.065 seeing 0.5"
03:45 ATLAS17fpq 16:00:06.21 +29:14:33.6 class 10min nothing is visible, aborted 
03:54 AT2017ehf  14:21:46.15 -01:23:15.2 class  20min Ia-norm at +17d z=0.055 seeing 0.7"
04:11 EG274      16:23:33.75 -39:13:47.5 15min 1.0" Gr11+Gr13+Gr16 seeing 0.7" very low S/N, useless 
04:30 AT2017ekg  17:04:10.30 +02:20:53.5 class 30min seeing 0.5" clouds low SNR
05:00 AT2017ekm  19:39:59.68 -01:48:19.9 class 30min seeing 0.6" clouds low SNR
05:35 AT2017ekl  20:14:42.06 -08:44:48.8 class 30min seeing 0.8" target clearly visible in the acquisition image, but nothing visible in the spectroscopic image, repeat the acquisition, nothing again, we decide to switch to SOFI
06:26 switch to SOFI
06:47 SN2017dfb  15:39:05.03 +05:34:16.9 30min JHK seeing 0.9"
07:16 LSQ15adm   20:14:21.82 -58:08:20.2 60min K seeing 0.9"
08:17 AT2016iks  23:07:21.52 +02:54:28.2 100min JHK seeing 0.8"
10:20 switch to EFOSC to perform a diagnostic test on the problem found at 05:35 UT, now everything seems to be OK, not clear what was the problem.
10:36 end of the observations
Comments